# AugInsert: Learning Robust Visual-Force Policies via Data Augmentation for Object Assembly Tasks

### v1

@@ V-A Training and Evaluation Details
In all experiments, we train the models for 100 epochs and perform 50 rollouts on the same task variations used in the training dataset. The model checkpoint that achieves the highest success rate in these rollouts during training is selected for evaluation. This training process is conducted over 6 random seeds per model, and the performance is averaged across all seeds during evaluation. A rollout is considered successful if it results in a successful insertion, and it is deemed failed if the maximum horizon is exceeded without insertion. Additionally, a rollout fails if the force-torque measurement surpasses a predefined threshold, to prevent unsafe behavior that could damage the robot arms or objects.
@@ V-B Evaluation on Unseen Task Variation Instances
To ensure rigorous evaluation, we explicitly separate task variation instances encountered during training from those used for evaluation within each variation category. For instance, when training with Grasp Pose variations, we include demonstrations involving x-axis translation and z-axis translation and rotation, while reserving y-axis rotation for evaluation. This approach guarantees that the model encounters unseen variations during evaluation, enabling us to assess its generalization to out-of-distribution inputs across all variation categories. A detailed overview of training versus evaluation instances for each variation category is provided in Table I.
@@
TABLE I: Instances for task variations during training (if included the training set) and evaluation. Task variations not in this table are the same both in training and evaluation.
@@ VI-A Which task variations are most difficult to generalize to?
Takeaways: We observe that Grasp Pose variations the hardest challenge to generalize to out of all of the individual task variations, as we see a drop from a mean success rate of 0.987 on Canonical rollouts with no task variations to 0.087 when Grasp Pose variations are applied. We hypothesize that the large negative impact on performance comes from the significant perturbation that grasp variations apply to all sensing modalities, unlike other variations such as Scene Appearance and Sensor Noise which only target specific modalities and thus have a smaller negative impact for the overall model.
@@ VI-B Which task variations included in the training set produce the largest impact on robustness?
To evaluate the effect of introducing task variations to the training dataset, we evaluate models trained on datasets augmented with different subsets of our task variations. These datasets contain the original collected human demonstrations as well as 6 augmentations per demonstration, with each augmentation containing a composition of all the task variations included in the dataset (which we refer to as the “training set variations” for that specific dataset). We evaluate these models both on instances of their training set variations that were unseen during training (as discussed in Section V-B) as well as all variations not included in the training set (which we refer to as the “evaluation set variations” for that specific dataset). In Figure 4 we report % success rate change from the canonical environment success rate averaged over 6 seeds for each variation, defined as
@@
Takeaways: We observe that the additions of Peg/Hole Shape, Object Body Shape, and Grasp Pose (considered the Base variations for this task) to the training set greatly reduce the generalization gap on unseen instances of these variations during evaluation. Curiously, we also observe a reduced generalization gap for a dataset with the Base training set variations on the evaluation set variations of Scene Appearance and Camera Pose, even though these variations had not been explicitly included in the dataset. This may be due to the similarity between the effects of applying grasp variations and perturbing the camera pose, as both alter the view of the object held by the gripper and the opposing object. Additionally, the resulting visual variations may have contributed to improving the model’s robustness to changes in scene appearance. Explicitly adding both visual variations (Scene Appearance and Camera Pose) and Sensor Noise to the training set does not further enhance generalization to their respective variations during evaluation.
@@ VI-C Can increasing the number of augmentations per demonstration improve robustness?
Takeaways: The most significant improvement in success rate as the number of augmentations increases seems to be the performance on Grasp Pose evaluations, with a more subtle upward trend in the other training set variations. Since the task variations of Peg/Hole Shape and Object Body Shape are discrete variations with a small subset of all possible shapes being included in the training set, they would benefit less from having more augmentations as the dataset would start to contain redundant instances of these variations. Grasp Pose variations, on the other hand, are continuous and so would benefit more from a larger sample of grasps. For the evaluation set variations (aside from All Variations), the success rate remains stable, suggesting that the model is not overfitting to the training set variations even when the dataset has more samples biased towards those perturbations.
@@ VI-D How much does each sensory modality contribute to model robustness?
Takeaways: We observe that out of the individual modalities, No Touch has the highest success rate drop for many of the variations (Peg/Hole Shape and Sensor Noise being the only exceptions) for both human-only and augmented demonstrations. On the other hand, No Vision has comparable (or sometimes even improved) % success rate changes to Full Model, suggesting its reduced significance in our overall framework compared to the other modalities. Since our task begins immediately in a contact state that is maintained throughout a majority of the task’s duration, it follows that force-torque provides the most valuable information about the task state. Proprioception may also give important task state information (as evidenced in the % success rate change for the No Prop. model), as the position of the two end-effectors relative to each other is highly correlated to the position of the peg and hole relative to each other, which is essential knowledge in completing the insertion task. Thus, visual observations provide the least relevant information for our task while still being susceptible to many of the task variations. However, visual input may still be essential in task contexts outside of the one studied here, especially in situations with little to no force-torque feedback (such as aligning the peg and hole objects to be in the same orientation before contact as was studied in our previous work [16]).
VI-E Additional Experiments
We perform a set of additional experiments to gain more insight into our system, the full results of which can be found in our supplementary material and website.
Attention Visualization: We plot attention weights in the cross-attention step of the visuotactile encoder averaged over the learned latent vectors and find that tokens from the tactile input take up a much larger proportion of attention than the visual tokens throughout the task, despite there being twice as many visual tokens as tactile ones. This provides further evidence of the importance of tactile over visual information first discussed in Section VI-D.
Real World Experiments: We construct a real-world setup of our insertion task and conduct a smaller-scale robustness experiment to compare with our results gathered in simulation.
VII Conclusion
We present a pipeline for data collection, augmentation, policy training, and evaluation to learn robust policies for an object assembly task across diverse observation-level task variations. Our experiments reveal that grasp variations pose the greatest challenge for generalization, and incorporating them through data augmentation significantly improves performance on unseen variations. Additionally, we demonstrate that force-torque input is critical for robustness to task variations, while the removal of RGB input has minimal impact.
APPENDIX
VII-A Task Variation Details
In this section, we provide more explicit details on the formulation and implementation of our task variations in simulation.
  1. 1.
    Peg and Hole Shape: The 9 possible peg and hole shapes are key, cross, diamond, line, circle, pentagon, hexagon, key, and u. The intrusions are built with a 5mm uniform tolerance on all sides.
  2. 2.
    Object Body Shape: The 3 possible object body shapes are cube (7.6cm on all sides), cylinder (7.6cm height and diameter), and octagonal prism (7.6cm height and diameter). The thin versions of these objects for the peg shape have $60\%$ the width of the full-sized shapes.
  3. 3.
[…]
  4. 4.
    Scene Appearance: Lighting variations involve both lighting color and the status of the light (either on or off). Object colors are sampled uniformly from all possible colors in the Robosuite [20] simulation, and floor textures were sampled from a discrete set of 20 possible textures (6 for training and 14 for evaluation).
  5. 5.
    Camera Pose: Camera position is perturbed in the range of [-4cm, 4cm] on the x, y, and z axes, and camera rotation is perturbed by sampling a random axis-angle rotation (with the axis sampled from a normalized spherical Gaussian and the angle sampled from [0°$\degree$, 10°$\degree$]) and applying it to the camera’s base orientation.
  6. 6.
    Sensor Noise: For force-torque readings, zero-mean Gaussian noise is applied to forces and torques separately (with a standard deviation of 5N for force and 0.15N-m for torque). For proprioceptive readings, zero-mean Gaussian noise is applied to end-effector position and orientation separately (with a standard deviation of 0.1cm for position and 0.57°$\degree$ for orientation).
VII-B Model Architecture
Figure 7: An extended view of our network architecture with added component dimensions and a visualization of our visuotactile transformer encoder.
In this section, we provide further details for our observation encoder and policy network architectures, an extended version of which can be found in Figure 7.
To encode our observations, we draw upon the success of visuotactile transformer encoders [2] and utilize a similar attention-based mechanism for RGB and tactile modality fusion. Rather than performing self-attention directly with the input tokens, we found that introducing a cross-attention step similar to the PerceiverIO [18, 19] architecture seemed to work best for our task. The PerceiverIO module has been used in other robotic manipulation contexts [26] to encode high-dimensional data while still maintaining a high-quality representation.
We tokenize our inputs by computing linear projections of both visual patches [27] for RGB inputs and individual readings per timestep for the force-torque input, and then add modality-specific position encodings. We then cross-attend these input tokens with a set of 8 learned latent vectors that then travel through a series of self-attention layers before ultimately being compressed and projected (as in [2]) to an output latent embedding $z_{vt}\in\mathbb{R}^{288}$. We encode proprioception with a multilayer perceptron to get an output embedding $z_{p}\in\mathbb{R}^{32}$ and concatenate $z_{vt}$ and $z_{p}$ to get $z\in\mathbb{R}^{320}$, which is used as input to the policy network. The policy network is then a multilayer perceptron that outputs 3-dimensional actions $a\in\mathcal{A}$ that represent end-effector position deltas for one of the arms.
VII-C Visuotactile Attention Visualization
To gain further insight into the information being learned by our model, we visualize the attention weights in the latent vector cross-attention step of the transformer visuotactile encoder. For each modality, we plot attention weights as proportions of total attention to tokens in that specific modality averaged over the 8 learned latent vectors. Figure 8 shows an example visualization; more examples can be found on our website.
Figure 8: A visualization of attention weights for vision and tactile tokens in the latent cross-attention step of the visuotactile transformer encoder. These weights are visualized as heatmaps overlayed on left and right wristview images for visual attention, and bars for each timestep under the force reading for tactile attention. We also plot the proportion of total attention for each modality (visual and tactile) during the course of a rollout.
Takeaways: Despite our model taking in twice as many visual tokens (72 tokens, 36 per view) as tactile ones (32 tokens), we observe that tactile attention takes up almost the entire proportion of attention across the input (as seen in the right-most plot of Figure 8). This finding provides further evidence to the importance of tactile information over visual information as discussed in Section VI-D, where we found that removing visual information from our input had little impact on the robustness of our model. Furthermore, we observe that the visual attention is mostly focused on semantically insignficant parts of the scene, such as the gripper at the bottom of the view, suggesting that the model is not receiving much useful visual information.
VII-D Comparing Data Augmentation Methods
In an effort to evaluate the validity of the online augmentation method for increasing the robustness of our model, we construct a dataset of human-generated trajectories with an extended set of visual variations and sensor noise, attempting to emulate a baseline data augmentation method that applies augmentations independently to each sensory modality offline during training. More specifically, we generate a dataset with training set variations of Scene Appearance (including object color, floor texture, and lighting), Camera Pose, and Sensor Noise with 12 augmentations per demonstration, but rather than keep applied variations consistent through each augmented rollout, we apply new instances of Scene Appearance and Camera Pose variations in each step of the demonstration. We also multiply the force and torque history reading by a random constant (from 0.1 to 2.0) independently determined each frame, following a similar data augmentation strategy used in [3]. We denote this dataset as Expanded Visual+Noise.
We report % success rate change from the Canonical environment success rate on models trained on the Expanded Visual+Noise dataset and compare it with the training set models from our original experiment (namely Visual+Sensor Noise that does not apply new variation instances per frame and Base that includes Grasp Pose variations); results can be found in Figure 9.
Figure 9: % success rate changes on each task variation for models trained on different subsets of task variations, including an additional Extended Visual+Noise dataset. Error bars represent one standard deviation from the mean. Our extended dataset of Scene Appearance, Camera Angle, and Sensor Noise variations does not lead to any significant improvements on robustness to any of the task variations over the original Visual+Sensor Noise dataset, particularly on Grasp Pose variations, where the inclusion of Grasp Pose in the training dataset via online augmentation still provides the largest boost in generalization ability.
Takeaways: We observe that our dataset with an expanded set of augmentations independently applied to each sensory modality does not necessarily improve robustness on most task variations (save for Peg/Hole Shape) compared to the original Visual+Sensor Noise dataset that was less extensive in terms of data augmentation. Most crucially, we do not see a significant improvement on Grasp Pose variations, validating the effect of non-independent multisensory data augmentation via trajectory replay. Thus, we have shown that even extensive independent augmentation of our multisensory input may not be enough to deal with certain task variations involved in our contact-rich task.
VII-E Success Rates for Canonical Environment
For full transparency for our experiments that involve reporting the % success rate change from the Canonical environment, we explicitly report the success rates in the no-variation Canonical environment, which the % success rate change is based off of, for each trained model. Average success rates over 6 training seeds are reported in Figures 10 and 11. It is worth noting that the average % success rate change across the 6 training seeds was calculated by determining the % success rate change for each individual seed and then calculating the average over those values, rather than calculating the average success rate across the 6 seeds first and then determining the difference of those averages.
Figure 10: Success rates on the Canonical environment for models trained on different training set variations. Error bars represent one standard deviation from the mean. These success rates correspond to the results reported in Figures 4 and 9.
Figure 11: Success rates on the Canonical environment for models using different modality combinations as input, trained on both human-only demonstration data as well as demonstration data augmented with Grasp Pose, Peg/Hole Shape, and Object Body Shape variations (denoted as the Base training set variations). Error bars represent one standard deviation from the mean. These success rates correspond to the results reported in Figure 6. We especially note the performance instability of the No Vision model trained on the Base training set variations, providing context for its omission in Figure 6.
VII-F Real-World Experimental Setup
Our real-world task setup, shown in Figure 12, is built to mirror our simulation setup as closely as possible. We designate one arm to be compliant, applying a constant amount of force while the other arm moves according to the actions given to it by the policy. In contrast to policies trained in simulation, our real-world policies predict 2-dimensional delta actions in the axes perpendicular to the axis of insertion (rather than 3-dimensional actions that include the axis along the direction of insertion), in order to prevent potentially unsafe interactions that may occur as a result of a premature insertion attempt. Once the peg and hole are aligned, the compliant arm automatically moves its held object forward to complete the insertion.
Figure 12: Our real-world experimental setup. Peg and hole object models are 3D-printed directly from the model files used in our simulation environment.
We train our real-world models with the same hyperparameters as those in simulation, although we only initiate 1 training seed per model (rather than 6). Additionally, we evaluate each model at the end of the entire training process, rather than performing training set rollouts during the training process to determine the best checkpoint. Success rates are determined by the number of successful trials over 20 rollouts. Successes and failures follow the same general criteria as in simulation, though a human manually annotates successes and failures per trial.
VII-G Real-World Experiments and Results
@@
Task Variation Difficulty: As a real-world analog to our experiments in Section VI-A, we evaluate a real-world policy trained on a dataset of human-generated demonstrations with no applied task variations on real-world versions of a subset of our task variations. Reported success rates over 20 rollouts can be found in Figure 13.
@@
Takeaways: Like in simulation, we observe that Grasp Pose variations seem to be the most difficult to generalize to, while the model is able to handle the mostly unisensory perturbations of Object Body Shape and Scene Appearance (object color and lighting). We also notice that our model struggles with Grasp Pose even when rotational grasp variations are removed; we hypothesize that this may be because a translational offset disrupts the desired behavior of lining up end-effector positions (given from proprioceptive input) in order to line up the peg and hole (i.e. solving the task can no longer be done by just matching the end-effector positions of the two arms). From these results, We believe that including Grasp Pose variations into the training dataset (as was done in simulation through online augmentation) may also improve performance in the real world.
@@
Takeaways: Like in simulation, we observe that the removal of force-torque data as input (the No Touch model) leads to a significant drop in success rate for all task variations compared to the Full Model, including the no-variations Canonical environment. We also see a small drop in performance for the No Vision model, somewhat aligning with our findings in simulation of the insignificance of visual input for our task. Surprisingly, we see performance increases in all task variations for the No Prop. model. We hypothesize that the small ranges of possible end-effector poses in our training dataset due to the high precision required for our task may cause our models to not learn much useful information from the proprioceptive embedding, though this observation may of course also be the result of a low sample size of trained models. Averaging the performance of models trained over multiple seeds (as was done in simulation), which was not able to be performed due to time constraints, may give us some more robust results.

### v2

@@ V-A Training and Evaluation Details
In all experiments, we train the models for 25000 training steps total. In simulation, we draw inspiration from Robomimic [25] and perform 50 rollouts every 1250 training steps on the same task variations used in the training dataset. The model checkpoint that achieves the highest success rate in these rollouts during training is selected for evaluation on unseen scenarios. This training process is conducted over 6 random seeds per model, and the performance is averaged across all seeds during evaluation. A rollout is considered successful if it results in a successful insertion, and it is deemed failed if the maximum horizon is exceeded without insertion. Additionally, a rollout fails if the force-torque measurement surpasses a predefined threshold (100N of force or 6N-m of torque in simulation), to prevent unsafe behavior that could damage the robot arms or objects.
We train our real-world models with the same hyperparameters as those in simulation, although we only initiate 1 training seed per model (rather than 6). Additionally, we evaluate each model at the end of the entire training process, rather than performing rollouts during the training process to determine the best checkpoint. Success rates are determined by the number of successful trials over 20 rollouts. Successes and failures follow the same general criteria as in simulation, though a human manually annotates successes and failures per trial.
@@ V-B Evaluation on Unseen Task Variation Instances
To ensure rigorous evaluation, we explicitly separate task variation instances encountered during training from those used for evaluation within each variation category. For instance, when training with Grasp Pose variations in simulation, we include demonstrations involving x-axis translation and z-axis translation and rotation, while reserving y-axis rotation for evaluation. This approach guarantees that the model encounters unseen variations during evaluation, enabling us to assess its generalization to out-of-distribution inputs across all variation categories. A detailed overview of training versus evaluation instances for each variation category is provided in Table I.
@@
TABLE I: Instances for task variations during training (if included in the training set) and evaluation. Task variations not in this table are the same both in training and evaluation.
@@ VI-A Which task variations are most difficult to generalize to?
Takeaways: We observe that Grasp Pose variations pose the greatest challenge for generalization out of all of the individual task variations, as we see a drop from a mean success rate of 0.987 on Canonical rollouts with no task variations to 0.087 when Grasp Pose variations are applied. We hypothesize that the large negative impact on performance comes from the significant perturbation that grasp variations apply to all sensing modalities, unlike other variations such as Scene Appearance and Sensor Noise which only target specific modalities and thus have a smaller negative impact for the overall model.
@@ VI-B Which task variations included in the training set produce the largest impact on robustness?
To evaluate the effect of introducing task variations to the training dataset, we evaluate models trained on datasets augmented with different subsets of our task variations. These datasets contain the original collected human demonstrations as well as 6 augmentations per demonstration, with each augmentation containing a composition of a subset of task variations (which we refer to as the “training set variations” for that specific dataset). We evaluate these models both on instances of their training set variations that were unseen during training (as discussed in Section V-B) as well as all variations not included in the training set (which we refer to as the “evaluation set variations” for that specific dataset). In Figure 5 we report % success rate change from the canonical environment success rate averaged over 6 seeds for each variation. Here, we define % success rate change as
@@
Takeaways: We observe that including only unisensory variations such as Scene Appearance (visual variations) and Sensor Noise do not significantly improve policy robustness with respect to multisensory variations such as Grasp Pose, signaling that unisensory data augmentation cannot handle multisensory perturbations. Although there is a limited amount of visual variations in this dataset, supplemental experiments showed that performance was not improved even with a greatly expanded set of visual and sensor noise variations. On the other hand, we observe that the additions of Peg/Hole Shape, Object Body Shape, and Grasp Pose (considered the Base variations for this task) to the training set greatly reduce the generalization gap on unseen instances of these variations during evaluation. Curiously, we also observe a reduced generalization gap for a dataset with the Base training set variations on the evaluation set variations of Scene Appearance and Camera Pose, even though these variations had not been explicitly included in the dataset. This may be due to the similarity between the effects of applying grasp variations and perturbing the camera pose, as both alter the view of the object held by the gripper and the opposing object. Additionally, the resulting visual variations may have contributed to improving the model’s robustness to changes in scene appearance. Explicitly adding both visual variations (Scene Appearance and Camera Pose) and Sensor Noise to the training set does not further enhance generalization to their respective variations during evaluation.
@@ VI-C Can increasing the number of augmentations per demonstration improve robustness?
Takeaways: The most significant improvement in success rate as the number of augmentations increases seems to be in the performance on Grasp Pose evaluations, with a more subtle upward trend in the other training set variations. Since the task variations of Peg/Hole Shape and Object Body Shape are discrete variations with a small subset of all possible shapes being included in the training set, they would benefit less from having more augmentations, as the dataset would start to contain redundant instances of these variations. Grasp Pose variations, on the other hand, are continuous and so would benefit more from a larger sample of grasps. For the evaluation set variations (aside from All Variations), the success rate remains stable, suggesting that the model is not overfitting to the training set variations even when the dataset has more samples biased towards those perturbations.
@@ VI-D How much does each sensory modality contribute to model robustness?
Takeaways: We observe that out of the individual modalities, No Touch has the highest success rate drop for many of the variations (Peg/Hole Shape and Sensor Noise being the only exceptions) for both human-only and augmented demonstrations. On the other hand, No Vision has comparable (or sometimes even improved) % success rate changes to Full Model, suggesting its reduced significance in our overall framework compared to the other modalities. Since our task begins immediately in a contact state that is maintained throughout a majority of the task’s duration, it follows that F/T data provides the most valuable information about the task state. Proprioception may also give important task state information (as evidenced in the % success rate change for the No Prop. model), as the position of the two end-effectors relative to each other is highly correlated to the position of the peg and hole relative to each other, which is essential knowledge in completing the insertion task. Thus, visual observations provide the least relevant information for our task while still being susceptible to many of the task variations. However, visual input may still be essential in task contexts outside of the one studied here, especially in situations with little to no F/T feedback (such as aligning the peg and hole objects to be in the same orientation before contact as was studied in our previous work [24]).
VI-E Are these empirical trends also shown in the real world?
@@
As a real-world analog to our experiments in Section VI-A, we evaluate a real-world policy trained on a dataset of 50 human-generated demonstrations with no applied task variations on real-world versions of a subset of our task variations. Reported success rates over 20 rollouts can be found in Figure 8.
@@
Takeaways: Like in simulation, we observe that Grasp Pose variations seem to be the most difficult to generalize to, while the model is able to handle the mostly unisensory perturbations of Object Body Shape and Scene Appearance (object color and lighting). We also notice that our model struggles with Grasp Pose even when rotational grasp variations are removed; we hypothesize that this may be because a translational offset disrupts the desired behavior of lining up end-effector positions given from proprioceptive input in order to line up the peg and hole (i.e. solving the task can no longer be done by just matching the end-effector positions of the two arms). From these results, we believe that including Grasp Pose variations into the training dataset (as was done in simulation through online augmentation) may also improve performance in the real world.
@@
Takeaways: Like in simulation, we observe that the removal of force-torque data as input (the No Touch model) leads to a significant drop in success rate for all task variations compared to the Full Model, including the no-variations Canonical environment. We also see a small drop in performance for the No Vision model, somewhat aligning with our findings in simulation of the insignificance of visual input for our task. Surprisingly, we see performance increases in all task variations for the No Prop. model. We hypothesize that the small ranges of possible end-effector poses in our training dataset due to the high precision required for our task may cause our models to not learn much useful information from the proprioceptive embedding, though this observation may of course also be the result of a low sample size of trained models. Averaging the performance of models trained over multiple seeds (as was done in simulation) may give us some more robust results.
VII Conclusion
Towards the goal of evaluating robustness of multisensory policies in contact-rich scenarios such as peg-in-hole object assembly, we present a pipeline for data collection, augmentation, policy training, and evaluation for an object assembly task across diverse observation-level task variations. Our experiments reveal that grasp variations pose the greatest challenge for generalization both in simulation and the real world, and incorporating them through data augmentation significantly improves performance on unseen variations. Additionally, we demonstrate that force-torque input is critical for robustness to task variations, while the removal of RGB input has minimal impact. For future work, we hope to expand our real-world experiments and validate online data augmentation as a method for increasing robustness in more realistic scenarios.